Could someone please help me on two logarithm word problems?
1)Recall that A = Ao2(-t/h). If the half-life of a substance is 10 years, how much of a 400 - gram sample will remain in 20 years?
2)Recall that A = Ao2 (-t/h). If the half-life of a substance is 4 days, how much of a 320-gram sample will remain in 6 days?
Thank you in advance for all the time all of you put into helping people like me.
Sincerely,
Elle J.

Answer by stanbon(48535)

1)Recall that A(t) = Ao2(-t/h). If the half-life of a substance is 10 years, how much of a 400 - gram sample will remain in 20 years?
----------
A(20) = 400* 2^(-20/10)
A(20) = 400* 2^-2
A(20) = 400* 1/4
A(20) = 100 grams
===============
2)Recall that A = Ao2 (-t/h). If the half-life of a substance is 4 days, how much of a 320-gram sample will remain in 6 days?
-----------------
A(6) = 320 *2^(-6/4)
A(6) = 320 *2^(-3/2)
A(6) = 320 *(0.35355...)
A(6) = 113.137... grams
=======================
Cheers,
Stan H.